# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

### v3

@@ 4.4.2. Enabling Multifaceted Self-Presentation Across Spaces and Relationships
Participants envisioned SMH supporting flexible, multifaceted self-presentation, making it easier to share different sides across various spaces and relationships while reducing pressure to maintain a single, polished identity. Rather than being limited to one polished identity, as on MSM, participants wanted ways to represent their full range of interests, identities, and relationships. P01 described this as having a house with rooms “dedicated to each of your interests,” explaining, “I’m into multiple things. You could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” This ability to represent different facets of oneself was seen as a way to reduce misjudgment and give others a more complete picture of who they are. P02 built on this idea, imagining that users with multiple “Muggle personalities” could organize them into separate spaces, saying, “Each personality has its own vault.”
@@ 4.4.3. Sharing Everyday Activities and Lived Experiences to Deepen Understanding and Connection
Participants valued sharing casual, everyday moments and rich, embodied memories, filling gaps between MSM’s typical highlights to create a more natural, meaningful extension of daily life and personal connection. P04 described wanting to share small, ordinary moments, like “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
Participants imagined SMH supporting personal, embodied memory-sharing to deepen understanding and reduce misinterpretation. P04 described wanting to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pen, [n. d.]) as a way to create stronger emotional connections. Unlike MSM, where communication often feels like “just words,” (P03) participants saw memory-sharing as a way to communicate context, emotion, and intention more fully. P02 explained that this kind of sharing could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
Some participants imagined these shared memories including sensory or emotional details beyond visuals. P01, for example, expressed interest in being able to “transmit smells through social media,” adding another layer to digital communication. Others compared it to live streaming personal experiences. P07 described it as “screen sharing, but with your eyes,” and P08 highlighted its usefulness in capturing fleeting moments, like crossing a state line, without the need to pause and record. P03 imagined being able to instantly share moments with loved ones “in the moment,” rather than having to explain them after the fact, when the emotion or relevance might already have passed.
@@ 4.4.4. Flexible, Contextual Privacy as a Foundation for Authentic Self-Presentation
Participants emphasized that flexible, contextual privacy is essential for authentic self-presentation, allowing people to manage who sees what, when, and how without feeling overexposed or misrepresented. They described current platforms as offering rigid and limited controls that fail to account for the complexity of real social life. P16 explained that on Instagram, for example, privacy options are reduced to features like “close friends,” adding, “it’s kind of a pain to organize.” P11 similarly noted that as people share “more like their genuine real life and their experiences,” privacy needs to become “a little bit more complicated,” requiring finer-grained control over visibility.
Participants also reflected on the challenge of managing different audiences and social contexts on MSM and imagined SMH supporting more seamless transitions between identities and spaces. P02 described a system that would “just telepathically know what’s in Muggle area and what’s in Hogwarts area,” allowing users to “walk through a portal” and appear as “a new person” without needing separate accounts. Building on this, participants imagined contextual identity management, where users could shift how they present themselves based on the audience or situation. P03 proposed a model where everyone maintains a “base” identity but customizes additional layers. P22 similarly described having different avatars or profiles for different groups, like on Discord, where “any group you’re in… you could change your avatar for that.”
Participants also stressed that contextual privacy should extend beyond people-based access to include temporal and relational factors. P14 proposed limiting access to different stages of one’s past identity, saying, “I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” Others, like P21, suggested gradual openness, where people could selectively reveal more of themselves over time. As they described, “if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on. Or turn it off.”
Participants also highlighted the importance of individual control over these transitions and boundaries. P14 explained that privacy settings should not be one-size-fits-all, suggesting that “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” Together, these suggestions reflect flexible, contextual privacy as not just a technical feature, but as a core requirement for supporting the kinds of personal, meaningful, and low-pressure interactions youth hoped SMH would make possible.
4.5. Intentional Engagement
Participants described growing frustration with current social media platforms prioritizing passive consumption over meaningful connection. In response, they envisioned SMH as a space promoting intentional, socially rewarding interactions. This section explores themes reflecting their desire to refocus social media on authentic connection, purposeful engagement, and meaningful time use.
@@ 4.5.1. Encouraging Deliberate Navigation and Content Discovery
Participants imagined SMH encouraging deliberate, effortful discovery, replacing MSM’s passive, algorithm-driven feeds with experiences requiring users to actively choose what and who they engage with. P22 described frustration with Instagram, explaining “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.” Instead, participants hoped SMH would encourage intentional, purposeful engagement when browsing content and navigating the virtual world.
P04 described how slowing down to explore someone’s space or story could make interactions feel more personal and meaningful. Unlike Instagram, where you “see all the posts at once” and can quickly scroll past them, SMH would require users to “take more time to hone in on a memory,” offering the satisfaction of consciously consuming content from someone you’re close to, rather than just “taking a quick second to look at it… and then being done.”
Participants imagined features prompting users to actively choose what they wanted to engage with before entering a space. P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces that would surface content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
Physical movement through SMH’s world was key to this intentionality. P03 explained that this kind of embodied movement makes SMH feel “more connected than regular social media,” (P03) because it requires active participation rather than passive scrolling. P02 preferred slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions. I want everything to be very thought through.” While convenient features existed, participants valued walking and exploring, turning simple actions into shared social moments.
P05 also imagined SMH as a place where users would intentionally seek out content or people, rather than having information pushed to them. They described SMH as a world where you might “walk through a town to get there,” explaining, “It’s not challenging to navigate, but you have to intentionally go there to see it. It isn’t something you can do mindlessly.” P14 agreed, noting that visiting someone’s space or content in SMH would feel much more personal than encountering “a random 2D Insta post that they posted 3 months ago.”
P05 believed this effortful navigation would naturally discourage superficial engagement: “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
@@ 4.5.2. Making Social Media About (Meaningful) Connection, Not (Passive) Consumption
Beyond content navigation, participants emphasized shifting the platform’s core purpose from endless consumption toward meaningful social connection. They expressed frustration with current platforms promoting mindless, passive consumption and hoped SMH would foster meaningful, intentional connections through active social engagement. They described frustration with how easily time slips away on platforms like Instagram and TikTok, leaving them unfulfilled. P14 captured this sentiment, describing spending time on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
@@
In contrast, participants described active, social engagement as far more rewarding. P16 shared that “anytime not scrolling” but “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 noted meaningful use comes when “actually talking to someone,” especially staying connected with people “not in close proximity.” They reflected that while users might spend more time on SMH, it would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to watching “50-second TikTok videos of somebody making slime for like an hour.”
@@
Reflecting on SMH’s potential, P09 acknowledged “time spent might not be mitigated,” but emphasized “the way you’re spending it” would make the difference. They explained SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content,” suggesting meaningful engagement, rather than passive consumption, ultimately mattered.
4.6. Playfulness and Whimsy
@@ 4.6.3. Augmenting Everyday Interactions with Fantastical Interfaces and Environments
4.7. Expanding Horizons
@@ 5. Discussion
Figure 2. 3D Gamer Youth’s Vision of Ideal Social Media Interactions
[…]

### v4

@@ 4.4.2. Enabling Multifaceted Self-Presentation Across Spaces and Relationships
SMH would support flexible, multifaceted self-presentation, making it easier to share different sides across various spaces and relationships without pressure to maintain a single, polished identity. P01 described having a house with rooms “dedicated to each of your interests”: “I’m into multiple things. You could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” Representing different facets would reduce misjudgment and give others a more complete picture. P02 imagined users with multiple “Muggle personalities” organizing them into separate spaces: “Each personality has its own vault.”
@@ 4.4.3. Sharing Everyday Activities and Lived Experiences to Deepen Understanding and Connection
Participants valued sharing casual, everyday moments and rich, embodied memories—filling gaps between MSM’s typical highlights to create a more natural extension of daily life. P04 wanted to share small, ordinary moments, like “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
SMH could also support personal, embodied memory-sharing to deepen understanding and reduce misinterpretation. P04 wanted to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pensieve). Unlike MSM, where communication often feels like “just words,” (P03) memory-sharing could communicate context, emotion, and intention more fully. P02 explained this could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
Some imagined shared memories including sensory or emotional details beyond visuals. P01 expressed interest in being able to “transmit smells through social media.” Others compared it to live streaming personal experiences—P07 described it as “screen sharing, but with your eyes,” and P08 highlighted its usefulness for capturing fleeting moments, like crossing a state line, without pausing to record. P03 imagined instantly sharing moments with loved ones “in the moment,” rather than explaining them after the fact when the emotion has passed.
@@ 4.4.4. Flexible, Contextual Privacy as a Foundation for Authentic Self-Presentation
Flexible, contextual privacy is essential for authentic self-presentation—allowing people to manage who sees what, when, and how without feeling overexposed or misrepresented. Current platforms offer rigid, limited controls that fail to account for real social life’s complexity. P16 explained that on Instagram, privacy options are reduced to features like “close friends”, adding, “it’s kind of a pain to organize.” P11 noted that as people share “more like their genuine real life and their experiences,” privacy needs become “a little bit more complicated,” requiring finer-grained visibility control.
SMH could support more seamless transitions between identities and spaces. P02 described a system that would “just telepathically know what’s in Muggle area and what’s in Hogwarts area,” allowing users to “walk through a portal” and appear as “a new person” without needing separate accounts. Users could shift how they present themselves based on audience or situation. P03 proposed a model where everyone maintains a “base” identity but customizes additional layers. P22 described having different avatars for different groups, like on Discord, where “any group you’re in… you could change your avatar for that.”
Contextual privacy would extend beyond people-based access to include temporal and relational factors. P14 proposed limiting access to different stages of one’s past identity: “I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” P21 suggested gradual openness, where people could selectively reveal more over time: “if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on. Or turn it off.”
Individual control over these transitions mattered. P14 explained that privacy settings should not be one-size-fits-all: “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” Together, these suggestions reflect flexible, contextual privacy not just as a technical feature, but as a core requirement for the personal, meaningful interactions youth hoped SMH would enable.
4.5. Theme 5: Intentional Engagement
Participants described growing frustration with current social media prioritizing passive consumption over meaningful connection. SMH would promote intentional, socially rewarding interactions—refocusing social media on authentic connection, purposeful engagement, and meaningful time use.
@@ 4.5.1. Encouraging Deliberate Navigation and Content Discovery
SMH would encourage deliberate, effortful discovery, replacing MSM’s passive, algorithm-driven feeds with experiences requiring users to actively choose what and who they engage with. P22 described frustration with Instagram: “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.”
P04 described how slowing down to explore someone’s space or story could make interactions feel more personal. Unlike Instagram, where you “see all the posts at once” and can quickly scroll past them, SMH would require users to “take more time to hone in on a memory,” offering the satisfaction of consciously consuming content from someone you’re close to, rather than “taking a quick second to look at it… and then being done.”
Some imagined features prompting users to actively choose what they wanted before entering a space. P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces surfacing content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
Physical movement through SMH’s world was key to this intentionality. P03 explained that embodied movement makes SMH feel “more connected than regular social media” because it requires active participation rather than passive scrolling. P02 preferred slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions. I want everything to be very thought through.”
P05 imagined SMH as a world where you might “walk through a town to get there”: “It’s not challenging to navigate, but you have to intentionally go there to see it. It isn’t something you can do mindlessly.” P14 agreed that visiting someone’s space in SMH would feel much more personal than encountering “a random 2D Insta post that they posted 3 months ago.”
This effortful navigation would naturally discourage superficial engagement. P05 noted: “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
@@ 4.5.2. Making Social Media About (Meaningful) Connection, Not (Passive) Consumption
Beyond content navigation, participants emphasized shifting the platform’s core purpose from endless consumption toward meaningful social connection. They expressed frustration with how easily time slips away on platforms like Instagram and TikTok, leaving them unfulfilled. P14 captured this, describing time spent on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
@@
Active, social engagement felt far more rewarding. P16 shared that “anytime not scrolling” but “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 noted meaningful use comes when “actually talking to someone,” especially staying connected with people “not in close proximity.” They reflected that while users might spend more time on SMH, it would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to watching “50-second TikTok videos of somebody making slime for like an hour.”
@@
P09 acknowledged “time spent might not be mitigated,” but emphasized “the way you’re spending it” would make the difference. SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content”—meaningful engagement, rather than passive consumption, ultimately mattered.
4.6. Theme 6: Playfulness and Whimsy
@@ 4.6.3. Augmenting Everyday Interactions with Fantastical Interfaces and Environments
4.7. Synthesis: SMH as a Hopeful Reframing of Social Media
@@ 5. Discussion
This paper starts from the premise that the central challenge in reimagining social media is not a lack of ideas, but a lack of conditions that allow different ideas to emerge. Asking young people how to redesign “social media” is itself a loaded prompt: the term carries accumulated controversy, cynicism, and tightly constrained expectations about what platforms can and should be. As our findings show, this framing activates design fixation rather than imagination. FI offers an alternative starting point by suspending the assumptions that quietly govern what feels reasonable, feasible, or even worth proposing. FI does not magically produce ideas that have never existed. Instead, it makes different configurations of priorities and interaction logic available to think with, and legible in participants’ own terms.
5.1. Defamiliarization, Not Novelty: How FI Expands the Frontier of Social Media Design
A useful way to interpret participants’ concepts is through defamiliarization: making strange what has been naturalized so we can notice the assumptions embedded in today’s dominant interaction grammar. Mainstream social media has converged on a paradigm organized around mobile-first interfaces, algorithmically curated feeds, frictionless consumption, and quantifiable performance. Under this convergence, many alternative logics are no longer legible as futures, even if their ingredients have existed before. Customizable profiles, spatial exploration, interest-based rooms, ambient co-presence, and playful identity work are not novel components. What is striking is how systematically they have been treated as side paths rather than viable organizing principles, crowded out by the priorities that dominate at scale.
“Newness” here should not mean inventing unprecedented features. Many elements participants reach for are historically familiar: MySpace made profile customization central; Second Life organized sociality through spatial exploration; chat rooms and MUDs normalized interest-based rooms and ambient co-presence. The problem is not that these ingredients never existed, but that they have been systematically crowded out by a particular configuration that won: the mobile-optimized, algorithmically-curated, vertically-scrolling feed oriented toward frictionless consumption and legible performance. The analytic question is less “has anyone ever built X?” than “why did X lose, and what does that loss reveal about what has become hard to imagine as a viable organizing principle?” When prior platforms are invoked today, they are often treated as feature catalogs or nostalgic references, rather than as prompts to reconsider the interaction logics and priorities that once made those logics coherent.
Seen this way, FI is not a machine for generating ideas no one has ever had; it is a method for making alternative configurations thinkable again. Even if every individual component has appeared before, specific arrangements matter enormously: the smartphone, touchscreen, and app store all existed before the iPhone, but a different bundling of affordances and defaults made a new paradigm visible. Participants’ proposals are revealing because they reorder emphasis: what is made reversible versus permanent, what requires consent, what counts as presence, and what the system treats as the default social unit. This reframing also helps explain why many youth examples draw from spaces not labeled “social media” at all: Discord servers, Minecraft worlds, and game-like environments where connection is organized around shared activity, rooms, and mutual orientation rather than broadcast audiences. FI loosens the category boundary itself, expanding what futures feel available to propose, including trajectories previously abandoned for reasons of incentive structures, scale narratives, or technical constraints rather than because their social ideas were exhausted.
FI matters because it interrupts the subtle constraints that ordinarily shape what participants feel permitted to propose. The fictional frame suspends economic, technical, and reputational pressures, but more importantly it removes the category label “social media,” which often triggers cynicism and incremental repair. In its place, FI invites reasoning from felt social experience: what presence should feel like, what kinds of disclosure should be low-stakes, what boundaries should be socially intelligible, and what kinds of connection should be effortless versus effortful. That shift makes it easier to surface configuration-level commitments: not whether a platform has a “profile” or “messaging,” but what counts as visibility, what is reversible, what requires consent, and what the system treats as the default social unit (an audience, a network, a room, a neighborhood, a party).
Figure 1. Youth’s Understanding of Social Media Interactions